# Text2Data: Low-Resource Data Generation with Textual Control

### v1

@@ Abstract
Natural language serves as a common and straightforward control signal for humans to interact seamlessly with machines. Recognizing the importance of this interface, the machine learning community is investing considerable effort in generating data that is semantically coherent with textual instructions. While strides have been made in text-to-data generation spanning image editing, audio synthesis, video creation, and beyond, low-resource areas characterized by expensive annotations or complex data structures, such as molecules, motion dynamics, and time series, often lack textual labels. This deficiency impedes supervised learning, thereby constraining the application of advanced generative models for text-to-data tasks. In response to these challenges in the low-resource scenario, we propose Text2Data, a novel approach that utilizes unlabeled data to understand the underlying data distribution through an unsupervised diffusion model. Subsequently, it undergoes controllable finetuning via a novel constraint optimization-based learning objective that ensures controllability and effectively counteracts catastrophic forgetting. Comprehensive experiments demonstrate that Text2Data is able to achieve enhanced performance regarding controllability across various modalities, including molecules, motions and time series, when compared to existing baselines. The code can be found at https://github.com/salesforce/text2data.
Machine Learning, ICML
Salesforce Research, 181 Lytton Ave, Palo Alto, CA 94301, USA
1 Introduction
@@
Autonomy and controllability stand as twin primary pillars of generative AI (Gozalo-Brizuela & Garrido-Merchan, 2023; Wang et al., 2022). While the challenge of autonomy has been substantially addressed through the rapid advancements of generative models, controllability is now ascending as a fervently explored arena within the machine learning community. As natural languages are one of the most common and simplest control signal for human beings to interact with machines, the machine learning community has increasingly focused on generating data that aligns semantically with textual descriptions, given its wide-ranging applications such as image editing (Zhang et al., 2023; Kawar et al., 2023), audio synthesis (Liu et al., 2023; Huang et al., 2023), video generation (Li et al., 2018; Hu et al., 2022), and many more (Tevet et al., 2023; Sanghi et al., 2022).
@@
Recent breakthroughs in text-to-data generative models, particularly those using diffusion techniques (Li et al., 2023; Yang et al., 2023; Kumari et al., 2023), have demonstrated remarkable proficiency by harnessing the rich semantic insights from vast datasets of data-text pairs. Despite the broad applicability of text-to-data generative models, not all modalities can meet the substantial data-text pair requirements for achieving optimal controllability during model training. This is often due to costly annotations or intricate data structures, a scenario we refer to as the low-resource situation. The lack of text labels in certain areas, such as molecules (Ramakrishnan et al., 2014; Irwin et al., 2012), motions (Guo et al., 2020; Mahmood et al., 2019), and time series (Du et al., 2020), primarily restricts supervised learning and hinders the use of advanced generative models for text-to-data generation tasks. The low-resource situation when training generative models unsurprisingly results in issues like undesirable generation quality, model overfitting, bias, and lack of diversity. As a result, the optimization for scarce text representations to improve the alignment between generated data and input texts in generative models is still under-explored.
@@
To mitigate the issues in the low-resource scenario, the strategies such as data augmentation (Hedderich et al., 2020; Meng et al., 2021), semi-supervised learning (Thomas et al., 2013; Cheuk et al., 2021), and transfer learning (Tits et al., 2020; Yi et al., 2018) are utilized. Yet, each comes across challenges. Data augmentation, for example, cannot always replicate genuine data fidelity to align accurately with initial text descriptions, and potentially leads to overfitting due to over-reliance on augmented samples. This method also exacerbates the training complexity, intensifying the already high computational demands of diffusion models. For semi-supervised learning, text inherently carries nuances, ambiguities, and multiple meanings. Ensuring that the model infers the correct interpretation when leveraging unlabeled data is not straightforward. Lastly, while transfer learning offers a solution for limited datasets, it is prone to catastrophic forgetting (Iman et al., 2023), where previous knowledge diminishes as new information (i.e., text descriptions) is introduced.
@@
  • We theoretically validate our optimization constraint selection and generalization bounds for our learning objective.
@@
2 Related works
2.1 Text-to-data diffusion-based generation
@@
Diffusion models, notably divided into classifier-guided (Dhariwal & Nichol, 2021) and classifier-free (Ho & Salimans, 2022) categories, have significantly impacted data generation across various domains. (Hoogeboom et al., 2022; Yang et al., 2023; Ho et al., 2022; Voleti et al., 2022). The classifier-guided diffusion guides the model during inference phase by independently training a classifier and supervising the model with its gradient, which is inefficient when computing gradient at each time step and sometimes the generation quality is deficient as the guidance is not involved in the training. By contrast, classifier-free diffusion guidance blends score estimates from both a conditional diffusion model and an unconditional one with time step as a parameter, exemplified by E(3) Equivariant Diffusion Model (EDM) (Hoogeboom et al., 2022) and Motion Diffusion Model (MDM) (Tevet et al., 2023) for controllable molecule and motion generation, respectively. Furthermore, since natural languages are a prevalent medium for human to communicate with the world, the text-to-data generation paradigm has gained traction, with diffusion models being instrumental in generating high-quality data aligned with textual inputs. The extensive applications encompass text-to-image generation (Ruiz et al., 2023; Zhang & Agrawala, 2023), text-to-speech generation (Huang et al., 2022; Kim et al., 2022), text-to-shape generation (Li et al., 2023; Lin et al., 2023), and more, leveraging the abundant text descriptions for training potent generative models. Despite advancements in generating data from text across various modalities, many other modalities may not satisfy the stringent requirements for sufficient data-text pairs essential for attaining optimal controllability during the training of models.
2.2 Low-resource learning
@@
In response to the challenges of low-resource training for controllable generative models, several strategies have been formulated. For instance, Yin et al. (2023) proposes Text-to-Text-to-Image Data Augmentation that employed both large-scale pretrained Text-to-Text and Text-to-Image generative models for data augmentation to generate photo-realistic labeled images in a controllable manner. Zang & Wan (2019) utilizes a semi-supervised approach to augment the training of both the encoder and decoder with both labeled and unlabeled data. Tu et al. (2019) proposes to learn a mapping between source and target linguistic symbols and employs transfer learning to transfer knowledge from a high-resource language to low-resource language. Nevertheless, all those strategies have their own limitations, such as high computational complexity for data augmentation, difficulty in maintaining the correct interpretation of text when leveraging unlabeled data during semi-supervised learning, and the potential catastrophic forgetting issues in transfer learning. Additionally, most of works on text-to-data generation lie in the modality of image, speeches, and texts, which have plenty of labeled datasets (Ito & Johnson, 2017; Pratap et al., 2020; Lin et al., 2014; Jiang et al., 2021; Wang et al., 2020) to train the models nowadays. Yet it is far under-explored for the low-source modalities such as molecules, motions and time series. Therefore, we propose Text2Data, a diffusion-based framework adept at harnessing limited text-labeled data to enhance the controllability of the model in text-to-data generation.
3 Problem formulation
@@ 4 Methods
where $p_{{\mathcal{D}}_{p}}({\mathbf{c}})$ is the true underlying text generating distribution corresponding to ${\mathbf{x}}\in{\mathcal{D}}_{p}$. Hence, as per Figure 1, we are inspired to initially utilize the unlabeled data in ${\mathcal{D}}$ to learn $p_{\theta}({\mathbf{x}})$ and obtain the optimal set of model parameters $\hat{\theta}\in\Theta$, which serves as a robust approximation of $p_{\theta}({\mathbf{x}}|{\mathbf{c}})$. Then, we finetune it using the data-text pairs in ${\mathcal{D}}_{p}$ to achieve desired model controllability (Section 4.1). Crucially, we ensure that the parameters remain in close proximity to those established when learning $p_{\theta}({\mathbf{x}})$ via constraint optimization to make Eq. 2 hold while mitigating catastrophic forgetting. Figure 1 shows our constraint to keep the finetuned parameters $\hat{\theta}^{\prime}$ within $\Theta\cap\Theta^{\prime}$, where $\Theta^{\prime}$ represents the optimal parameters from unconstrained finetuning. Then we plug the generative diffusion implementation to our framework in Section 4.2. Finally, we use Section 4.3 to offer theoretical insights into the selection of the optimization constraint, underpinned by generalization bounds for our learning objective.
4.1 Learning text-to-data generation under low-resource situation
@@
where $p_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}})$ is the true underlying data-text joint distribution. $\Theta$ denotes a localized parameter space where a minimum can be located. Specifically, we minimize $\mathbb{E}_{{\mathbf{x}},{\mathbf{c}}\sim p_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}})}[-\log p_{\theta}({\mathbf{x}}|{\mathbf{c}})]$ using the labeled data in ${\mathcal{D}}_{p}$ within the optimal set $\{\theta:\mathbb{E}_{{\mathbf{x}}\sim p_{{\mathcal{D}}_{p}}({\mathbf{x}})}[-\log p_{\theta}({\mathbf{x}})]\leq\xi\}$ to make the parameter not far from learned via Eq. 3, so that catastrophic forgetting is mitigated. The trade-off between the learning objective and its constraint aligns with the nature of a lexicographic optimization problem (Gong & Liu, 2021) and therefore can be solved in that context.
4.2 Generative objective on empirical samples
Eq. 4 is the population-level learning objective while empirically we follow the standard loss function (i.e., transformed evidence lower bound of Eq. 4) of the classifier-free diffusion guidance (Ho & Salimans, 2022) by optimizing:
@@ 4.2 Generative objective on empirical samples
where we have: ${\mathcal{L}}_{1}(\theta)=\mathbb{E}_{{\mathbf{x}}\sim p_{{\mathcal{D}}}({\mathbf{x}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)-{\bm{\epsilon}}||^{2}]$, ${\mathcal{L}}_{1}^{\prime}(\theta)=\mathbb{E}_{{\mathbf{x}}\sim p_{{\mathcal{D}}_{p}}({\mathbf{x}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)-{\bm{\epsilon}}||^{2}]$ and ${\mathcal{L}}_{2}(\theta)=\mathbb{E}_{{\mathbf{x}},{\mathbf{c}}\sim p_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},{\mathbf{c}},t)-{\bm{\epsilon}}||^{2}]$. Also, $t$ is sampled from uniform between 1 and $T$, $T$ is the total diffusion steps, ${\bm{\epsilon}}$ is the standard Gaussian random variable, and ${\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)$ and ${\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},{\mathbf{c}},t)$ are functions we aim to fit at the $t$-th diffusion step. Note that ${\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)$ and ${\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},{\mathbf{c}},t)$ share the same parameters but are just trained at different stages: distribution mastery on unlabeled data and controllable finetuning on labeled data, respectively. The framework of classifier-free diffusion model and the derivation of ${\mathcal{L}}_{1}(\theta)$, ${\mathcal{L}}_{1}^{\prime}(\theta)$ and ${\mathcal{L}}_{2}(\theta)$ are introduced in Appendix A.
@@
where we have ^$\hat{{\mathcal{L}}}_{1}(\theta)=\mathbb{E}_{{\mathbf{x}}\sim\hat{p}_{{\mathcal{D}}}({\mathbf{x}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)-{\bm{\epsilon}}||^{2}]$, ^$\hat{{\mathcal{L}}}_{1}^{\prime}(\theta)=\mathbb{E}_{{\mathbf{x}}\sim\hat{p}_{{\mathcal{D}}_{p}}({\mathbf{x}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)-{\bm{\epsilon}}||^{2}]$ and ^$\hat{{\mathcal{L}}}_{2}(\theta)=\mathbb{E}_{{\mathbf{x}},{\mathbf{c}}\sim\hat{p}_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},{\mathbf{c}},t)-{\bm{\epsilon}}||^{2}]$. $\hat{p}_{{\mathcal{D}}}({\mathbf{x}})$, $\hat{p}_{{\mathcal{D}}_{p}}({\mathbf{x}})$ and $\hat{p}_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}})$ are corresponding empirical data generating distributions. $\hat{\Theta}$ is the localized parameter space where a minimum can be located for ^$\hat{{\mathcal{L}}}_{1}(\theta)$. The lexicographic optimization of Eq. 6 is presented in Algorithm 1, which relies on learning the update direction on model parameter (i.e., ^$\nabla\hat{\mathcal{L}}_{2}(\theta)+\lambda\nabla\hat{{\mathcal{L}}}_{1}^{\prime}(\theta)$ in Algorithm 1) to balance the trade-off between the minimization of ^$\hat{{\mathcal{L}}}_{2}(\theta)$ and ^$\hat{{\mathcal{L}}}_{1}^{\prime}(\theta)$ using a dynamic gradient descent (Gong & Liu, 2021):
@@
where $\alpha$, $\beta$ and $\gamma$ are predefined positive hyperparameters. More details and intuition of lexicographic optimization are explained in Appendix B.

### v2

@@ Abstract
Natural language serves as a common and straightforward signal for humans to interact seamlessly with machines. Recognizing the importance of this interface, the machine learning community is investing considerable effort in generating data that is semantically coherent with textual instructions. While strides have been made in text-to-data generation spanning image editing, audio synthesis, video creation, and beyond, low-resource areas characterized by expensive annotations or complex data structures, such as molecules, motion dynamics, and time series, often lack textual labels. This deficiency impedes supervised learning, thereby constraining the application of advanced generative models for text-to-data tasks. In response to these challenges in the low-resource scenario, we propose Text2Data, a novel approach that utilizes unlabeled data to understand the underlying data distribution through an unsupervised diffusion model. Subsequently, it undergoes controllable finetuning via a novel constraint optimization-based learning objective that ensures controllability and effectively counteracts catastrophic forgetting. Comprehensive experiments demonstrate that Text2Data is able to achieve enhanced performance regarding controllability across various modalities, including molecules, motions and time series, when compared to existing baselines. Code—https://github.com/SalesforceAIResearch/text2data Extended version—https://arxiv.org/abs/2402.10941
Introduction
@@
Autonomy and controllability stand as twin primary pillars of generative AI (Gozalo-Brizuela and Garrido-Merchan 2023; Wang et al. 2022). While the challenge of autonomy has been substantially addressed through the rapid advancements of generative models, controllability is now ascending as a fervently explored arena within the machine learning community. As natural languages are one of the most common and simplest control signal for human beings to interact with machines, the machine learning community has increasingly focused on generating data that aligns semantically with textual descriptions, given its wide-ranging applications such as image editing (Zhang et al. 2023; Kawar et al. 2023), audio synthesis (Liu et al. 2023; Huang et al. 2023), video generation (Li et al. 2018; Hu, Luo, and Chen 2022), and many more (Tevet et al. 2023; Sanghi et al. 2022).
@@
Recent breakthroughs in text-to-data generative models, particularly those using diffusion  (Li et al. 2023; Yang et al. 2023; Kumari et al. 2023), have demonstrated remarkable proficiency by harnessing the rich semantic insights from vast datasets of data-text pairs. Despite the broad application of text-to-data generative models, not all modalities can meet the substantial data-text pair requirements for achieving optimal controllability during model training. This is often due to costly annotations or intricate data structures, a scenario we refer to as the low-resource situation. The lack of text labels in certain areas, such as molecules (Ramakrishnan et al. 2014; Irwin et al. 2012), motions (Guo et al. 2020; Mahmood et al. 2019), and time series (Du et al. 2020), primarily restricts supervised learning and hinders the use of advanced generative models for text-to-data generation tasks. The low-resource situation when training generative models unsurprisingly results in issues like undesirable generation quality, model overfitting, bias, and lack of diversity. However, the optimization for scarce text representations to improve the alignment between generated data and input texts in generative models is still under-explored.
@@
To mitigate the issues in the low-resource scenario, strategies such as data augmentation (Hedderich et al. 2020; Meng et al. 2021), semi-supervised learning (Thomas et al. 2013; Cheuk, Herremans, and Su 2021), and transfer learning (Tits, El Haddad, and Dutoit 2020; Yi et al. 2018) are utilized. Yet, each comes across challenges. Data augmentation, for example, cannot always replicate genuine data fidelity to align accurately with initial text descriptions, and potentially leads to overfitting due to over-reliance on augmented samples. It also exacerbates the training complexity, intensifying the already high computational demand of diffusion models. For semi-supervised learning, text inherently carries nuances, ambiguities, and multiple meanings. Ensuring that the model infers the correct interpretation when leveraging unlabeled data is not straightforward. Lastly, while transfer learning offers a solution for limited datasets, it is prone to catastrophic forgetting (Iman, Arabnia, and Rasheed 2023), where previous knowledge diminishes as new text descriptions are introduced.
@@
  • We theoretically validate optimization constraint selection and generalization bounds for our learning objective.
@@
Related works
Text-to-data diffusion-based generation
@@
Diffusion models, notably divided into classifier-guided (Dhariwal and Nichol 2021) and classifier-free (Ho and Salimans 2022) categories, have significantly impacted data generation across various domains. (Hoogeboom et al. 2022; Yang et al. 2023; Ho et al. 2022; Voleti, Jolicoeur-Martineau, and Pal 2022). The classifier-guided diffusion guides the model during inference phase by independently training a classifier and supervising the model with its gradient, which is inefficient when computing gradient at each time step and sometimes the generation quality is deficient as the guidance is not involved in the training. By contrast, classifier-free diffusion guidance blends score estimates from both a conditional diffusion model and an unconditional one with time step as a parameter, exemplified by E(3) Equivariant Diffusion Model (EDM) (Hoogeboom et al. 2022) and Motion Diffusion Model (MDM) (Tevet et al. 2023) for controllable molecule and motion generation, respectively. Furthermore, since natural languages are a prevalent medium for human to communicate with the world, the text-to-data generation paradigm has gained traction, with diffusion models being instrumental in generating high-quality data aligned with textual inputs. The extensive applications encompass text-to-image generation (Ruiz et al. 2023; Zhang and Agrawala 2023), text-to-speech generation (Huang et al. 2022; Kim, Kim, and Yoon 2022), text-to-shape generation (Li et al. 2023; Lin et al. 2023), and more, leveraging the abundant text descriptions for training potent generative models. Despite advancements in generating data from text across various modalities, many other modalities may not satisfy the stringent requirements for sufficient data-text pairs essential for attaining optimal controllability during the training of models.
Low-resource learning
@@
In response to the challenges of low-resource training for controllable generative models, several strategies have been formulated. For instance, Yin et al. (2023) proposes Text-to-Text-to-Image Data Augmentation that employed both large-scale pretrained Text-to-Text and Text-to-Image generative models for data augmentation to generate photo-realistic labeled images in a controllable manner. Zang and Wan (2019) utilizes a semi-supervised approach to augment the training of both the encoder and decoder with both labeled and unlabeled data. Tu et al. (2019) proposes to learn a mapping between source and target linguistic symbols and employs transfer learning to transfer knowledge from a high-resource language to low-resource language. Nevertheless, all those strategies have their own limitations, such as high computational complexity for data augmentation, difficulty in maintaining the correct interpretation of text when leveraging unlabeled data during semi-supervised learning, and the potential catastrophic forgetting issues in transfer learning. Additionally, most of works on text-to-data generation lie in the modality of image, speeches, and texts, which have plenty of labeled datasets (Ito and Johnson 2017; Pratap et al. 2020; Lin et al. 2014; Jiang et al. 2021; Wang, Wu, and Pino 2020) to train the models nowadays. Yet it is far under-explored for the low-source modalities such as molecules, motions and time series. Therefore, we propose Text2Data, a diffusion-based framework adept at harnessing limited text-labeled data to enhance the controllability of the model in text-to-data generation.
Problem formulation
@@ Methods
where $p_{{\mathcal{D}}_{p}}({\mathbf{c}})$ is the true underlying text generating distribution corresponding to ${\mathbf{x}}\in{\mathcal{D}}_{p}$. Hence, as per Figure 1, we are inspired to initially utilize the unlabeled data in ${\mathcal{D}}$ to learn $p_{\theta}({\mathbf{x}})$ and obtain the optimal set of model parameters $\hat{\theta}\in\Theta$, which serves as a robust approximation of $p_{\theta}({\mathbf{x}}|{\mathbf{c}})$. Then, we finetune it using data-text pairs in ${\mathcal{D}}_{p}$ to achieve desired model control. Crucially, we ensure that the parameters remain in close proximity to those established when learning $p_{\theta}({\mathbf{x}})$ via constraint optimization to make Eq. 2 hold while mitigating catastrophic forgetting. Figure 1 shows our constraint to keep the finetuned parameters $\hat{\theta}^{\prime}$ within $\Theta\cap\Theta^{\prime}$, where $\Theta^{\prime}$ represents the optimal parameters from unconstrained finetuning. Then we plug the generative diffusion implementation to our framework. Finally, we use the last section to offer theoretical insights into the selection of the optimization constraint, underpinned by generalization bounds for our learning objective.
Learning text-to-data generation under low-resource situation
@@
where $p_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}})$ is the true underlying data-text joint distribution. $\Theta$ denotes a localized parameter space where a minimum can be located. Specifically, we minimize $\mathbb{E}_{{\mathbf{x}},{\mathbf{c}}\sim p_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}})}[-\log p_{\theta}({\mathbf{x}}|{\mathbf{c}})]$ using the labeled data in ${\mathcal{D}}_{p}$ within the optimal set $\{\theta:\mathbb{E}_{{\mathbf{x}}\sim p_{{\mathcal{D}}_{p}}({\mathbf{x}})}[-\log p_{\theta}({\mathbf{x}})]\leq\xi\}$ to make the parameter not far from learned via Eq. 3, so that catastrophic forgetting is mitigated. The trade-off between the learning objective and its constraint aligns with the nature of a lexicographic optimization problem (Gong and Liu 2021) and therefore can be solved in that context.
Generative objective on empirical samples
Eq. 4 is the population-level learning objective while empirically we follow the standard loss function (i.e., transformed evidence lower bound of Eq. 4) of the classifier-free diffusion guidance (Ho and Salimans 2022) by optimizing:
@@
where we have: ${\mathcal{L}}_{1}(\theta)=\mathbb{E}_{{\mathbf{x}}\sim p_{{\mathcal{D}}}({\mathbf{x}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)-{\bm{\epsilon}}||^{2}]$, ${\mathcal{L}}_{1}^{\prime}(\theta)=\mathbb{E}_{{\mathbf{x}}\sim p_{{\mathcal{D}}_{p}}({\mathbf{x}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)-{\bm{\epsilon}}||^{2}]$ and ${\mathcal{L}}_{2}(\theta)=\mathbb{E}_{{\mathbf{x}},{\mathbf{c}}\sim p_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},{\mathbf{c}},t)-{\bm{\epsilon}}||^{2}]$. Also, $t$ is sampled from uniform between 1 and $T$, $T$ is the total diffusion steps, ${\bm{\epsilon}}$ is the standard Gaussian random variable, and ${\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)$ and ${\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},{\mathbf{c}},t)$ are functions we aim to fit at the $t$-th diffusion step. Note that ${\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)$ and ${\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},{\mathbf{c}},t)$ share the same parameters but are just trained at different stages: distribution mastery on unlabeled data and controllable finetuning on labeled data, respectively. The framework of classifier-free diffusion model and the derivation of ${\mathcal{L}}_{1}(\theta)$, ${\mathcal{L}}_{1}^{\prime}(\theta)$ and ${\mathcal{L}}_{2}(\theta)$ are introduced in Appendix.
@@
where we have ^$\hat{{\mathcal{L}}}_{1}(\theta)=\mathbb{E}_{{\mathbf{x}}\sim\hat{p}_{{\mathcal{D}}}({\mathbf{x}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)-{\bm{\epsilon}}||^{2}]$, ^$\hat{{\mathcal{L}}}_{1}^{\prime}(\theta)=\mathbb{E}_{{\mathbf{x}}\sim\hat{p}_{{\mathcal{D}}_{p}}({\mathbf{x}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},t)-{\bm{\epsilon}}||^{2}]$ and ^$\hat{{\mathcal{L}}}_{2}(\theta)=\mathbb{E}_{{\mathbf{x}},{\mathbf{c}}\sim\hat{p}_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}}),t}[||{\bm{\epsilon}}_{\theta}({\mathbf{x}}^{(t)},{\mathbf{c}},t)-{\bm{\epsilon}}||^{2}]$. $\hat{p}_{{\mathcal{D}}}({\mathbf{x}})$, $\hat{p}_{{\mathcal{D}}_{p}}({\mathbf{x}})$ and $\hat{p}_{{\mathcal{D}}_{p}}({\mathbf{x}},{\mathbf{c}})$ are corresponding empirical data generating distributions. $\hat{\Theta}$ is the localized parameter space where a minimum can be located for ^$\hat{{\mathcal{L}}}_{1}(\theta)$. The lexicographic optimization of Eq. 6 is presented in Algorithm 1, which relies on learning the update direction on model parameter (i.e., ^$\nabla\hat{\mathcal{L}}_{2}(\theta)+\lambda\nabla\hat{{\mathcal{L}}}_{1}^{\prime}(\theta)$ in Algorithm 1) to balance the trade-off between the minimization of ^$\hat{{\mathcal{L}}}_{2}(\theta)$ and ^$\hat{{\mathcal{L}}}_{1}^{\prime}(\theta)$ using a dynamic gradient descent (Gong and Liu 2021):
@@
where $\alpha$, $\beta$ and $\gamma$ are predefined positive hyperparameters. More details and intuition of lexicographic optimization are explained in Appendix.